### v3

# Heuristic Adaptation of Potentially Misspecified Domain Support for Likelihood-Free Inference in Stochastic Dynamical Systems

Abstract
In robotics, likelihood-free inference (LFI) can provide the domain distribution that adapts a learnt agent in a parametric set of deployment conditions. LFI assumes an arbitrary support for sampling, which remains constant as the initial generic prior is iteratively refined to more descriptive posteriors. However, a potentially misspecified support can lead to suboptimal, yet falsely certain, posteriors. To address this issue, we propose three heuristic LFI variants: EDGE, MODE, and CENTRE. Each interprets the posterior mode shift over inference steps in its own way and, when integrated into an LFI step, adapts the support alongside posterior inference. We first expose the support misspecification issue and evaluate our heuristics using stochastic dynamical benchmarks. We then evaluate the impact of heuristic support adaptation on parameter inference and policy learning for a dynamic deformable linear object (DLO) manipulation task. Inference results in a finer length and stiffness classification for a parametric set of DLOs. When the resulting posteriors are used as domain distributions for sim-based policy learning, they lead to more robust object-centric agent performance.
I Introduction
Consider guiding a deformable linear object (DLO) using only vision and proprioception (visuomotor) near a stack of cubes, to remove (whip) the top cube from the stack (fig. 1, left). As high-dimensional visual states can be inefficient for parameter inference and policy learning and deployment, we use inferred keypoints [1, 2, 3] to track the DLO (whip) and cubes, thus having more efficient lower-dim states [4].
Robotics simulators are becoming increasingly accurate in representing advanced dynamics [5], such as DLO behaviour. They are used to collect large amounts of physical interaction data and train control policies for challenging tasks, such as dynamically manipulating a DLO [6, 7, 8]. Calibrating a simulator’s parameterisation $\boldsymbol{\theta}$ to the real-world (Real2Sim[9, 10] is crucial to close the reality gap, which is the discrepancy between the behaviour of a real system and its simulated twin. Likelihood-free inference (LFI) [11], such as BayesSim and subsequent works [2, 12, 13, 14, 15], models the probability density of a multimodal posterior $\hat{p}(\boldsymbol{\theta})$ as a mixture of Gaussians (MoG). The modes of this mixture are the main hypotheses of $\boldsymbol{\theta}$, with uncertainty modelled as the corresponding variances.
This probabilistic approach is integrable with domain randomisation (DR) [16], which is used to achieve effective sim-based policy training and robust policy deployment in the real world (Sim2Real[17, 8]. DR trains a reinforcement learning (RL) policy in simulated environments parameterised using samples drawn from a domain distribution (dd) [18]. These domain samples robustify the Sim2Real deployment towards their corresponding conditions. Thus, a wide and uniform dd leads to a generalist agent, whereas a narrow and concentrated dd leads to a specialist agent.
Figure 1: The support misspecification issue for a visuomotor DLO whipping task (left, timelapse). On the centre (red boxes), we see two cases of how LFI struggles to accurately infer the Young’s modulus and length posterior on a misspecified support, and the implications on domain samples drawn from the posterior. On the right (green boxes), we see how an adapted support leads to a more accurate inference result and more descriptive sets of domain samples. Orange arrows denote the accumulation of posterior density on domain boundaries, which signals potential misspecification. Blue arrows denote the corresponding adaptations, which stretch the domain.
This integrated Real2Sim2Real treatment requires the a priori definition of an initial prior for sampling in parameter inference or DR. This is commonly a uniform distribution with finite support in a range $\Theta=[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ [19, 20, 2, 12, 15]. However, in the absence of theoretical or practical evidence, intuitively defining this domain support is not always practical or possible and can lead to prior misspecification, which is a fundamental concern in Bayesian methods [21]. In particular, in sequential Bayesian inference [22] it is common for the support $\Theta$ to be constant, even as the prior is refined from an initial uniform distribution to more concentrated MoGs. This is particularly problematic in LFI for Real2Sim, where we inherently do not have intrinsic knowledge of the system.
To understand the implications of maintaining an arbitrary support $\Theta$ when addressing the reality gap of a simulation, let us go back to our DLO whipping task and try to hit the top cube, while avoiding any impact with the rest of the stack. Let the key physical parameters of the task be the whip’s length and stiffness. For stiffness, even knowing the source material may not be enough to accurately define $\Theta$. On the material side, the curation process that created the DLO may have altered the original mechanical properties [23]. On the simulator side, although robotics simulators have increasingly realistic physics, realism does not necessarily imply physical accuracy [24]. That is, the parameterisation needed to match the observed behaviour of a real object in simulation may differ from its actual physical properties.
From the above, we see not only that it is challenging to accurately define a prior support $\Theta$ in which an LFI algorithm will search for the target system parameterisation $\boldsymbol{\theta}$, but that it can also be undesirable since we risk excluding useful value regions while biassing the parameter inference, and consequently policy search, with a false assumption.
For example, what if the actual length or stiffness exists beyond $\Theta$? In fig. 1 (centre), we see the posterior mass accumulating at the boundaries of the given domain. This is how a misspecified $\Theta$ can lead to misguided estimates, since the LFI attributes $\mathbf{x}^{r}$ to the closest $\Theta$ region, even with low posterior $\hat{p}$ confidence. We will see that it is non-trivial to determine how low this $\hat{p}$ confidence, quantified as probability density, should be for $\hat{p}$ to be ineffective (section V-B). In addition, a misspecified $\Theta$ can affect policy learning, when the inferred posterior is used as a domain distribution.
One might ask “Why not use the widest support possible?”. For example, in the whipping task, let us widen $\Theta$ by setting its extrema to the lowest and highest possible length and stiffness that we can think of. This minimises the risk of ignoring useful parameterisations, and we can expect that the granularity of the samples will be progressively refined through inference iterations (albeit not as much as when sampling within a much narrower domain). We will see that such an approach is only partially useful, as it exposes us to reduced data efficiency and reduced data quality (section III-D).
For example, depending on the simulator’s physics solver configuration, there will be a lower softness threshold, simulating below which is infeasible, as it results in DLO mesh collapse, inducing internal collisions. This can lead to simulation failure [5], leading to failure of the entire learning process or to storing data without information value. Handling such cases online in data-intensive tasks, such as simulation-based inference and policy learning, is non-trivial [25, 26]. Directly tuning a simulator for more robustness, e.g. by increasing physics solver iterations, can significantly reduce computational efficiency, which is an undesirable trade-off. Thus, in tasks such as our DLO whipping, we define a physically feasible domain $\Phi\supset\Theta$ to restrict parameter search from exploring infeasible value regions.
Due to the above, we need LFI extensions that enable the adaptation of a misspecified support $\Theta$ over inference steps. Our key insight is that searching for the $\Theta$ boundaries that would improve the estimation of $\boldsymbol{\theta}$ intuitively resembles an information acquisition problem that searches for what data to collect next to maximise information gain on $\theta$. This is treated by Bayesian experimental design (BED) [27] and, specifically, Bayesian optimisation (BO) [28, 29].
BO iteratively refines where to sample in an input space using a surrogate model and an acquisition function. However, directly using BO for support adaptation would require the evaluation of possible $\Theta$ redefinitions, which is intractable, as each $\Theta$ candidate would require running a complete inference step of costly simulations. Still, we observe that the position and shape of a posterior hint at whether we should adapt the boundaries of $\Theta$ and in what direction (fig. 1, centre). Thus, if we consider that the evolving posterior $\hat{p}(\boldsymbol{\theta})$ proxies a surrogate model, we need to define a proxy acquisition function that will adapt $\Theta$ to progressively focus our sampling where most promising. We do this by heuristically evaluating properties such as $\hat{p}(\boldsymbol{\theta})$ mass cumulation or mode shift toward $\Theta$ boundaries.
I-A Contributions
First, we expose how support misspecification can lead to suboptimal posterior inference despite perceived certainty of the results in two stochastic dynamical benchmarks, the Lotka-Volterra and the M/G/1 queue model (section III).
Second, we propose (section IV) and validate over Lotka-Volterra and M/G/1 (section V) three BayesSim variants, each using a different support adaptation heuristic to guide support expansion over inference steps. Our heuristics are compatible with existing LFI algorithms [30, 31, 12], as they can directly extend them, using information that has already been computed in a typical Bayesian inference step.
Third, we integrate our most robust variant into a Real2Sim2Real framework for our DLO whipping task (section VI). Then we show that support adaptation refines the inferred physical properties of a DLO drawn from a parametric set. To our knowledge, we are the first to address the reality gap for a parameter space that combines extrinsic (length) and intrinsic (softness) material properties (section VII).
Fourth, we use the resulting posteriors, with and without support adaptation, for DR in RL policy learning in simulation and then show the real-world object-centric performance impact of these posteriors on our DLO task (section VII).
Using two low-dim benchmarks and a high-dim visuomotor DLO control task, we validate heuristic support adaptation across various stochastic dynamical systems. We show that for a set of (real) DLOs, adapting domain support during iterative LFI can refine physical property inference among visually similar objects. Through DR, this fine inference translates into a stronger object-centric specialisation of RL agents, with measurable real-world performance impact.
II Likelihood-free inference for adaptive domain randomisation
In this section, we cover the background on LFI (section II-A), iterative posterior refinement (section II-B), adaptive DR (section II-C, section II-D) and BED (section II-E), which is necessary to understand our heuristic support adaptation contribution (section IV).
II-A Likelihood-free inference
LFI treats a simulator as a black-box generative model $g$ with intractable likelihood, which uses its parameterisation $\boldsymbol{\theta}$ to generate an output $\mathbf{x}$ that represents the behaviour of the simulated system [11]. This process defines a likelihood function $p(\mathbf{x}\mid\boldsymbol{\theta})$, which cannot be evaluated, but can be indirectly sampled by running the simulator on a sampled $\boldsymbol{\theta}$. The intractability of the simulator’s likelihood is an important challenge for simulation-based inference [32] and has motivated its approximate Bayesian computation (ABC) [33] subfield, the precursor of contemporary LFI methods.
In robotics, overcoming the reality gap requires solving the inverse problem of mapping real observations $\mathbf{x}^{r}$ to the parameters $\boldsymbol{\theta}$ that are most likely to produce them in simulation; let $\mathbf{x}^{s}=g(\boldsymbol{\theta})$. This defines the problem of approximating the posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}^{s},\mathbf{x}^{r})$.
BayesSim [19] approximates the posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$ by learning the conditional density function (CDF) $q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x})$, parameterised by $\boldsymbol{\phi}$. The CDF is modelled as a mixture of Gaussians (MoG), which can be approximated by mixture density neural networks (MDNN) [34]. The inputs $\mathbf{x}$ can be high-dimensional state-action trajectories, summary statistics $\psi(\cdot)$ of trajectory data, or kernel mappings.
BayesSim first requires training $q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x})$ on a dataset of $N$ pairs $(\boldsymbol{\theta}_{n},\mathbf{x}_{n})$, where the parameters $\boldsymbol{\theta}_{n}$ are drawn from a proposal prior $\tilde{p}(\boldsymbol{\theta})$ and the observation trajectories $\mathbf{x}_{n}$ are generated by running $g(\boldsymbol{\theta}_{n})$ for the duration of a simulated episode and collecting the respective state-action pairs.
Then, given a single real-world trajectory $\mathbf{x}^{r}$, BayesSim estimates the posterior as:
$\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto\frac{p(\boldsymbol{\theta})}{\tilde{p}(\boldsymbol{\theta})}q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\text{,}$ (1)
which allows flexibility (likelihood-free) for a desirable prior $p(\boldsymbol{\theta})$ which is different from the proposal prior. If $\tilde{p}(\boldsymbol{\theta})=p(\boldsymbol{\theta})$, then $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$.
II-B Iterative posterior refinement
Equation 1 is used for sequential Bayesian inference, as shown in fig. 2 and alg. 1, with a colour coding that connects illustration and formulation. Sequential posterior refinement is performed by progressively adapting the sampling distribution to regions of higher posterior density. This is useful since sampling across a wide parameter space is inefficient when only a narrow region of $\boldsymbol{\theta}$ yields simulated observations similar to $\mathbf{x}^{r}$. This generalised LFI formulation has been used for Real2Sim inference [2] and dual control [31, 30]. In this paper, when referring to posterior refinement and Bayesian inference, we use the terms sequential and iterative interchangeably, since the latter can communicate our experiments with more technical precision.
In addition to iterative posterior refinement, sequential inference can involve progressive densification of the dataset $\mathcal{D}$ used to train the CDF $q_{\phi}$ (alg. 1, line 10[35]. In this way, our belief about $\boldsymbol{\theta}$ evolves recursively as more informative simulated evidence accumulates. When the samples drawn from each iteration’s posterior are aggregated in the same dataset, densifying certain areas of the supported parameter space, this proxies the reweighing of the prior to adaptively shift focus toward informative regions. Thus, in practice, we can ignore the proposal prior $\tilde{p}(\boldsymbol{\theta})$ and infer the posterior using only the (desirable) prior $p(\boldsymbol{\theta})$ and the CDF.
This dataset-oriented approach has the benefit of avoiding direct manipulation of the prior distribution, which can be numerically unstable. This resembles Bayesian adaptive inference [36] and aligns with the broader paradigm of Bayesian filtering [37, 38], in which a belief in latent quantities is updated recursively given accumulating evidence.
[…]
Figure 2: Sequential refinement of Bayesian posterior approximations. Each posterior $\hat{p}_{t}$ is used to sample new parameters, generate simulations, and retrain the density estimator. This adaptive loop densifies coverage in high-likelihood regions.
Algorithm 1 Iterative Bayesian LFI
1:function BayesSim($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r}$)
2:  Args: $N_{\text{LFI}}$: inference iterations;
3:     $\tilde{p}$: Proposal prior, usually $\tilde{p}(\boldsymbol{\theta})\approx\mathit{U}$;
4:     $\pi_{\boldsymbol{\beta}_{0}}$: Initial policy;
5:     $\mathbf{x}^{r}$: Target trajectory, collected with $\pi_{\boldsymbol{\beta}_{0}}$
6:  $p_{0}\leftarrow\tilde{p} \triangleright$ Init. reference prior
7:  $i\leftarrow 0;\;\mathcal{D}\leftarrow\{\}$ $\triangleright$ Init. samples dataset
8:  while $i<N_{\text{LFI}}$ do
9:   $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}\leftarrow$ Simulate NN $\pi_{\boldsymbol{\beta}_{0}}$ rollouts, $\boldsymbol{\theta}\sim p_{i}$
10:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}$ $\triangleright$ Dataset augmentation
11:   Train $q_{\phi}$ over $\mathcal{D}$
12:   $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto(p_{i}(\boldsymbol{\theta})\mathbin{/}\tilde{p}(\boldsymbol{\theta}))q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$
13:   $i\leftarrow i+1$
14:   $p_{i}\leftarrow\hat{p} \triangleright$ Update reference prior
15:  end while
16:  return $\hat{p} \triangleright$ Return latest posterior
17:end function
II-C Domain randomisation for reinforcement learning
Domain randomisation aims to provide enough simulated variability at training time so that at test time the model generalises to real-world data [16]. This happens by randomising the simulation’s parameterisation vector through samples drawn from a domain distribution. However, when the inherent instability of training RL agents is combined with DR using wide uniform priors, we may not get the robustness expected in certain tasks [39, 31]. With LFI, we obtain a posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$ that is qualitatively more precise, i.e. narrower than the uniform prior, and more accurate, i.e. dense around the system’s true parameters.
RL aims to maximise the expected sum of future discounted rewards (by $\gamma$) following a policy $\pi_{\boldsymbol{\beta}}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, parameterised by $\boldsymbol{\beta}$ and sampling action $\mathbf{a}_{t}$ given state $\mathbf{s}_{t}$. We formulate DR for RL as maximising the joint objective:
$\mathcal{J}(\boldsymbol{\beta})=\mathbb{E}_{\boldsymbol{\theta}}\Biggl[\mathbb{E}_{\boldsymbol{\eta}}\Biggl[\sum_{t=0}^{T-1}\gamma^{(t)}r(\mathbf{s}_{t},\mathbf{a}_{t})\mid\boldsymbol{\beta}\Biggr]\Biggr]\text{,}$ (2)
with respect to policy’s $\boldsymbol{\beta}$, where $\boldsymbol{\theta}\sim\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$ and $\boldsymbol{\eta}=\{\mathbf{s}_{t},\mathbf{a}_{t},\mathbf{o}_{t}\}_{t=0}^{T-1}$ the history of observation $\mathbf{o}_{t}$, action $\mathbf{a}_{t}$ pairs over time horizon $T$. In this paper, we consider $\mathbf{s}_{t}\equiv\mathbf{o}_{t}$.
II-D Adaptive domain randomisation
To our knowledge, we are the first to perform support adaptation for Real2Sim parameter inference; however, there is some DR work on adapting domain ranges to produce generalist agents, robust to environment perturbations for a given task [40, 41, 42]. Starting with a best estimate of the randomisation limits [43], these methods adapt the domain ranges, usually by expanding, during model training or RL policy learning. The adaptation guides the DR sampling towards parameterisations that challenge the current version of the model [25, 26, 44], thus robustifying it to a progressively broader scope of deployment conditions.
Such approaches hypothesise that, when trained on a very wide domain distribution, the model parameters act as an implicit memory mechanism, implementing a learning algorithm that adjusts agent behaviour to improve performance in the deployment environment over time [40]. However, this is infeasible in short-horizon dynamic tasks, such as DLO whipping [6, 45, 7], as they are too short (typically $10$\sec$\$ and <<32<<32 steps) for such a behaviour adaptation.
We start with a best estimate of the support [43], which we expand [40] when certain heuristic rules are met. However, we perform LFI and integrate into a Real2Sim2Real framework that efficiently and interpretably produces a specialist agent for a given real world environment [45, 14, 3]. Both efficiency and interpretability rest in training RL policies using inferred posteriors as domain distributions. In this way, our DR reflects our estimate on the parameterisation of the deployment environment and therefore induces object-centric agent behaviour [46].
II-E Bayesian experimental design & Bayesian optimisation
BED [27] searches for the next experiment to be run to improve our estimate of the parameter vector $\boldsymbol{\theta}$. It assumes a generative model $p(\mathbf{y}\mid\mathbf{x},\boldsymbol{\theta})$, where $\mathbf{x}$ is controllable experimental conditions, in our case the domain support, and $\mathbf{y}$ is the observed outcome. It uses a utility function $U(\mathbf{x};\mathbf{y},\boldsymbol{\theta})$ that scores how informative an experiment is. This function is normally expected information gain (EIG):
$U(\mathbf{x})=\mathbb{E}_{\mathbf{y}\sim p(\mathbf{y}\mid\mathbf{x})}[\textbf{KL}(p(\boldsymbol{\theta}\mid\mathbf{y},\mathbf{x})\mid\mid p(\boldsymbol{\theta})].$ (3)
BED has been used in adaptive design contexts, where we make a series of design decisions and use previous results at each step [21]. In this context, BO has emerged as a special case of adaptive, or sequential, BED [47, 48, 49, 50]. BO is used to optimise black-box, expensive to evaluate and noisy functions [28, 29]. Instead of directly searching the problem space, it builds a probabilistic surrogate model of the function $f$ and uses this model and an acquisition function (AF) to decide where to sample next.
The surrogate model is usually a Gaussian process (GP), which provides a mean prediction $\mu(\mathbf{x})$ and an uncertainty estimate $\sigma(\mathbf{x})$. The AF uses the surrogate model to balance exploration, i.e., sampling high-uncertainty regions, and exploitation, i.e., sampling regions with expectedly high $f$ values. Common AFs are probability of improvement, which evaluates $f$ at the point most likely to improve its value, expected improvement (EI), which evaluates $f$ where most likely to improve $f^{\prime}$ the most, thus avoiding getting stuck in local optima, and upper confidence bound, which explicitly decomplexes exploration and exploitation when selecting the next $f$ evaluation point.
III A demo of the support misspecification issue
Now that we have overviewed iterative posterior refinement and BED, we can inspect the impact of a misspecified support. We do this through the following questions.
  1. 1.
    What does the misspecified support issue look like?
  2. 2.
    Why not use the widest support possible?
We experiment with the stochastic dynamical benchmarks of the Lotka-Volterra and M/G/1 queue, as implemented for the seminal LFI work of [11]. The hyperparameters of the MDNN used to approximate the CDF (alg. 1, line 12) in the following experiments are given in App. B, Table II.
III-A Lotka-Volterra predator-prey population model
The Lotka-Volterra model describes the interaction dynamics between populations of two species, commonly a predator ($X$) and a prey ($Y$). It is based on differential equations that describe how the rise of one species’ population affects the other, leading to circles and fluctuations on the species sizes. In this context, there are four possible species interactions: 1. a predator is born, 2. a predator dies, 3. a prey is born, 4. a prey dies being eaten by a predator. The rate of each interaction is controlled by the respective positive parameter in a set $\boldsymbol{\theta}=\{\theta_{1},\theta_{2},\theta_{3},\theta_{4}\}$.
The Lotka-Volterra model can be simulated by evaluating a set of nonlinear ordinary differential equations (ODEs): $\frac{dX}{dt}=\theta_{1}XY-\theta_{2}X$ and $\frac{dY}{dt}=\theta_{3}Y-\theta_{4}XY$. However, it can also be formulated as a Markov jump process (MJP), in which predator and prey populations change via discrete, random events occurring in continuous time with rates derived from $\boldsymbol{\theta}$ [11]. Specifically, we draw the time to the next reaction from an exponential distribution with rate equal to the total rate $\theta_{1}XY+\theta_{2}X+\theta_{3}Y+\theta_{4}XY$. We then select a reaction $\in[1,4]$ at random, with probability proportional to its $\boldsymbol{\theta}$ rate, simulate it, and repeat. The MJP version of the model exhibits complex and oscillatory dynamics, similar to the ODE, but with randomness replacing the dependence on initial conditions [51]. In both cases, nature-like observations emerge only for specific $\boldsymbol{\theta}$ values, leading to narrow posteriors that are difficult to recover under intractable likelihood, making this model a popular ABC benchmark.
Figure 3: Population counts over $30$ timesteps of 44 sample Lotka-Volterra system simulations, recording with $dt=0.2$, for $(X,Y)=(50,100)$ and $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0.01,0.5,1.0,0.01)$, demonstrating stochastic variability in its oscillatory behaviour due to the MJP formulation.
For initial populations $(X,Y)=(50,100)$, a ground truth $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0.01,0.5,1.0,0.01)$ can produce the typical oscillatory behaviour of a realistic predator/prey populations system (fig. 3). For this $\boldsymbol{\theta}$, an OK support has been shown to be $5,2]\times 4 [$11]. Given this, a misspecified support for $\theta_{1}$ and $\theta_{2}$ would be [[3.,2.],[5,1.5],[5,2.],[5,2.]][[-3.,2.],[-5,-1.5],[-5,2.],[-5,2.]]. In addition, we empirically find that a broad support is $6,4]\times 4$, a broader support is $6,5]\times 4$, and the broadest support which gives a reasonable simulation success rate is $7,7]\times 4$. We conduct our Lotka-Volterra experiments assuming these parameter spaces, simulating for $30$ time units and recording $(X,Y)$ values with $dt=0.2$. Domain names are underlined following the plot and title background colours in fig. 5.
III-B M/G/1 queue model
The M/G/1 model describes how a server manages a queue of continuously arriving jobs. The job arrivals are Markovian (Poisson process), the service time, i.e. the time required for the server to process a job and remove it from the queue, follows a general distribution, and there is just $\mathbf{1}$ server, hence M/G/1. The service time is independently and uniformly distributed in a range $\theta_{1},\theta_{2}]$. The time between two consecutive job arrivals is independently and exponentially distributed with rate $\theta_{3}$. The server can only observe the time elapsed between the departure of two consecutive jobs. More formally, if $s_{i}$ the service time of job $i$, which enters the queue in time $u_{i}$ and leaves in time $d_{i}$, we have $s_{i}\sim\mathit{U}(\theta_{1},\theta_{2})$, $u_{i}-u_{i-1}\sim\exp(\theta_{3})$, and $d_{i}-d_{i-1}=s_{i}+\max(0,u_{i}-d_{i-1})$. Thus, to simulate a target trajectory, we need to infer $\boldsymbol{\theta}=\{\theta_{1},\theta_{2},\theta_{3}\}$.
The M/G/1 queue is a stochastic discrete event process and, like the Lotka–Volterra, is a popular ABC benchmark. Both models generate stochastic trajectories, but for different reasons: in Lotka–Volterra (MJP formulation), the randomness arises from demographic noise in birth–death interactions, while in M/G/1 it is driven by random arrival times and service durations. For M/G/1, the variability introduced by these processes can make the observed data $\mathbf{x}$ less directly informative about the parameters $\boldsymbol{\theta}$, thus increasing the difficulty of inference (fig. 4). In fact, assuming a total of $I=50$ jobs, we only know a set of percentiles of the empirical distribution of jobs’ interdeparture times (idts) $d_{i}-d_{i-1},i\in[1,I]$. Following [11], we use $5$ equally spaced percentiles of the set of idts: 0th, 25th, 50th, 75th and 100th.
Figure 4: Interdeparture time (idt) analysis for 33 sample M/G/1 system simulations for $50$ jobs, demonstrating stochastic variability in jobs’ completion.
We assume a ground-truth of $(\theta_{1},\theta_{2},\theta_{3})=(1.0,5.0,0.2)$, for which $[[0,10],[0,10],[0,0.35]]$ has been shown to be an OK support [11]. Given this, a misspecified support for $\theta_{1}$ and $\theta_{2}$ would be $[[3.0,10.0],[0.0,7.0],[0,0.35]]$. We also consider a broad support being $[[0,20],[0,20],[0,0.5]]$. The underline colours correspond to App. A, fig. 18.
III-C What does the misspecified support issue look like?
Figure 5: The misspecified support issue and the difficulty of compensating with a broader support for Lotka-Volterra. We plot the progression of prior samples (top) and their respective MoG posteriors (bottom) along $10$ inference iterations. On prior samples scatterplots, the heatmap indicates the iteration each sample was drawn, the lighter the colour, the later the iteration. Each iteration samples’ bounding box is plotted in dashed lines of the respective colour. The bigger dots mark the accumulated dataset’s mean in each inference iteration. A black dashed line shows this trajectory. On MoG heatmaps, MoG progression is annotated with arrows pointing to the position of the components’ weighted mean for each iteration and colorbars quantify likelihood. For coherence, we plot only the last iteration’s posterior.
Figure 5 shows how a wider support (3rd col.) not only does not guarantee stable convergence to ground truth (compared to the 1st col.), but instead hinders the learning of a reliable CDF among inference iterations. We also see how a misspecified support, which does not include the ground truth, leads to ineffective inference (2nd col.). Interestingly, we observe that the adversarial effects of misspecifying the $\theta_{1}$ and $\theta_{2}$ support also carry over to $\theta_{3}$ and $\theta_{4}$ inference, despite their subdomains being adequately specified.
We empirically find that Lotka-Volterra is a more helpful task in exposing the misspecified support issue, since it features configurations of $\boldsymbol{\theta}$ that can lead to infeasible simulations. Thus, the support is better implemented as a strict effective support for a distribution, by clipping outlier samples accordingly [^1] In scientific computing and deep learning libraries, such as NumPy and PyTorch, this corresponds to using their built-in .clip functions.. Modern robotics simulators often have similar physically infeasible configurations, which can be intractable to robustify against. The respective illustration and discussion for M/G/1 is available in App. A.
III-D Why not use the widest support possible?
Figure 6: The reduced data efficiency of sampling over various wide supports for $10$ LFI iterations for Lotka-Volterra. On the left, we plot the attempted simulations per iteration (top) and the number of failures (bottom). On the right, we plot the simulation success rate per iteration (top) and the cumulative number of failed simulations per iteration (bottom).
Figure 6 shows more implications of progressively widening the posterior. We see that for Lotka-Volterra, which inherently has infeasible parameter space configurations, sampling from a wider prior leads us to launching infeasible simulations that fail and have no information value. This is most evident in the bottom-right cumulative sim failures plot. In contrast, we see how sampling from progressively narrower posteriors, when the inference converges on a belief with increased certainty, leads to a higher simulation success rate. This emerges as a performance pattern, since we see that even for suboptimal broad and broader support experiments, the simulation success rate improves in later iterations. We understand that data efficiency is correlated with posterior sharpness, since both generally increase during iterations.
Overall, we see that the narrower the support, the more likely an efficient data collection is. A wider support can reduce the success rate, requiring more simulations to converge to a posterior that may be precise but inaccurate. All experiments were performed with $\max 125$ simulations (successful or failed) per iteration.
IV Heuristic support adaptation methods
Our main insight is that adapting a misspecified support resembles searching for the experiment configuration that would maximise the information on $\boldsymbol{\theta}$ in BED. However, integrating alg. 1 into BED would be intractable, as it would require computing the EIG over possible support redefinitions. Also, we will not be sampling directly in the input space of the simulator; instead, we will be adapting the domain support over inference iterations.
We present three heuristic extensions to alg. 1. Each heuristic answers “Given what I know about $\hat{p}(\boldsymbol{\theta}|\mathbf{x}=\mathbf{x}^{r})$, how should I adapt the domain support to increase the value of subsequent inference?” in its own way. One heuristic checks probability mass accumulation near the support bounds and expands accordingly (section IV-A). Another checks MoG mode shift towards the bounds in order to expand (section IV-B). Our third heuristic treats the weighted mean of the MoGs as the centre of the desired support (section IV-C).
We draw inspiration from BO’s use of tractable acquisition functions as surrogates for intractable EI objectives. However, while BO’s surrogate is a GP posterior over the unknown $\boldsymbol{\theta}$, our surrogate is the MoG posterior over $\boldsymbol{\theta}$. While BO’s acquisition maps $(\mu(\mathbf{x}),\sigma(\mathbf{x}))$ to a utility score to pick next $\mathbf{x}$, our heuristics compare map posterior features, such as mass accumulation, mode location, and boundary proximity, with adaptation thresholds, which, if exceeded, shape the next domain support as an expansion of the current.
Thus, our BayesSim variants reduce the support adaptation rule to a deterministic heuristic. Each heuristic acts as an information acquisition function, which does not optimise a formal expected utility, but encodes implicit value estimates that mimic the exploration/exploitation balance.
We treat support expansion as window stretching [40] rather than window sliding [9]: instead of shifting the sampling window, we iteratively accumulate new parameter samples guided by the latest posterior (alg. 1, line 10). We compactly denote a distribution $p$ with support $\Theta\subseteq\mathbb{R}^{(D)}$ as $p_{(\Theta)}$, but do this sparingly. This arbitrarily denotes that the distribution is supported by a $D$-dim. rectangular range defined by the bounds $\Theta=[\boldsymbol{\theta}_{\min}^{(d)},\boldsymbol{\theta}_{\max}^{(d)}]_{d=1}^{D}$.
IV-A BayesSim-EDGE: expanding bounds via edge mass
Algorithm 2 BayesSim-EDGE
1:function BayesSim_EDGE($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r},\Theta_{0},$...)
2:  Args: (for $N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r}$ see alg. 1);
3:     $\Theta_{0}=[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$: initial support bounds;
4:     $\delta$: edge zone fraction; $\tau$: prob. mass thresh;
5:     $\eta$: edge expansion factor;
6:     $\Phi$: phys. feasible domain limits
7:  $D\leftarrow\lvert\boldsymbol{\theta}_{\min}\rvert;\;p_{0}\leftarrow\tilde{p};\;i\leftarrow 0;\;\mathcal{D}\leftarrow\{\}$
8:  while $i<N_{\text{LFI}}$ do
9:   $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}\leftarrow$ Simulate NN $\pi_{\boldsymbol{\beta}_{0}}$ rollouts, $\boldsymbol{\theta}\sim p_{i}$
10:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}$ $\triangleright$ Dataset augmentation
11:   Train $q_{\phi}$ over $\mathcal{D}$
12:   $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto(p_{i}(\boldsymbol{\theta})\mathbin{/}\tilde{p}(\boldsymbol{\theta}))q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$
13:   // EDGE: Check and adapt parameter bounds
14:   for $d=1$ to $D$ do
15:     $r^{(d)}=\lvert\theta_{\max}^{(d)}-\theta_{\min}^{(d)}\rvert \triangleright$ Range size
16:     $\Delta_{L}^{(d)}\leftarrow[\theta_{\min}^{(d)},\theta_{\min}^{(d)}+\delta]$ $\triangleright$ Left edge zone
17:     $\Delta_{R}^{(d)}\leftarrow[\theta_{\max}^{(d)}+\delta,\theta_{\max}^{(d)}]$ $\triangleright$ Right edge zone
18:     // Accum. edge masses
19:     Accum. $M_{L}^{(d)}$ in $\Delta_{L}^{(d)}$ & $M_{R}^{(d)}$ in $\Delta_{R}^{(d)}$
20:     if $M_{L}^{(d)}>\tau\And(\theta_{\min}^{(d)}-\eta r^{(d)})\in\Phi$ then
21:      $\theta_{\min}^{(d)}\leftarrow\theta_{\min}^{(d)}-\eta r^{(d)}$ $\triangleright$ Expand left
22:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\min}^{(d)}\leftarrow\Phi_{\min}^{(d)}$
23:     if $M_{R}^{(d)}>\tau\And(\theta_{\max}^{(d)}+\eta r^{(d)})\in\Phi$ then
24:      $\theta_{\max}^{(d)}\leftarrow\theta_{\max}^{(d)}+\eta r^{(d)}$ $\triangleright$ Expand right
25:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\max}^{(d)}\leftarrow\Phi_{\max}^{(d)}$
26:   end for
27:   $i\leftarrow i+1$
28:   $\Theta_{i}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ $\triangleright$ Update bounds
29:   $p_{i}\leftarrow\hat{p}_{(\Theta_{i})}$ $\triangleright$ Update reference prior
30:  end while
31:  return $\hat{p}$
32:end function
BayesSim-EDGE (Edge-Driven Gaussian Expansion) expands the support bounds based on the accumulation of posterior mass near them (alg. 2). The motivation is to expand the parameter search space when the learnt posterior exhibits a significant mass close to existing bounds, suggesting that the true parameters may lie beyond them.
We begin with standard BayesSim assumptions, meaning a proposal prior $\tilde{p}$ from which we sample the parameters $\boldsymbol{\theta}$ and a policy $\pi_{\boldsymbol{\beta}_{0}}$ to be used for data collection both in sim and in real (for $\mathbf{x}^{r}$). We assume the proposal prior to be uniform and we expand our formulation to concretely define the initial support as $\Theta_{0}=[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$, or, compactly, $\tilde{p}_{(\Theta_{0})}$.
EDGE is configurable through its hyperparameters. These are the edge mass threshold $\tau$, the edge zone fraction $\delta$, the edge expansion factor $\eta$, and the physically feasible domain limits $\Phi$. Specifically, $\delta$ is the size of the area within which we consider any accumulated probability mass to be near the edge. This is the sensitivity of our expansion criterion, which compares the accumulated probability mass with a threshold $\tau$. We denote the left and right edge zones ($\min$ and $\max$ side) as $\Delta_{L}^{(d)}$ and $\Delta_{R}^{(d)}$ for a dimension $d$ of $\boldsymbol{\theta}$, and the respective accumulated masses as $M_{L}^{(d)}$ and $M_{R}^{(d)}$. If $\tau$ has been exceeded and if an expansion would still be within $\Phi$ (lines 20, LABEL: and 23), we consider the expansion criterion to be satisfied and expand the respective domain boundary by a factor of $\eta$ (lines 21, LABEL: and 24). If $\tau$ has been exceeded, but the expansion would have been beyond the extremum of the respective $\Phi$ limit, then this becomes the new boundary.
We denote the hyperparameters as scalars, but they can also be vectors, e.g. $\tau\rightarrow\boldsymbol{\tau}$. This allows customising heuristic rules for each posterior dimension, e.g. having a different edge mass threshold for the $\hat{p}(\boldsymbol{\theta})$ dimension referring to a DLO Young’s modulus and a different one for its length. This holds for all heuristic hyperparameters going forward.
EDGE explores beyond the initial parameter ranges if the inferred posterior (line 12) confidently points toward the edges of the domain. This confidence is tuned through $\tau$ and $\delta$. Thus, within $N_{\text{LFI}}$ iterations, we can converge to an adapted support $\Theta_{N_{\text{LFI}}}$ in conjunction with a posterior $\hat{p}$.
Conceptually, EDGE dictates that if there is an accumulation of $\hat{p}$ mass near an edge, then we should extend the support in this direction. This rule treats posterior mass accumulation as a geometric indication that there is more value in exploring beyond the corresponding $\Theta$ boundary. In this process, $\tau$ and $\delta$ balance exploration/exploitation. With a broad $\delta$ and a low $\tau$, EDGE favours exploration, easily expanding in the direction of the probability mass. With a narrow $\delta$ and a large $\tau$, EDGE favours exploitation of the current $\Theta$, exploring beyond its limits only when given a strong signal towards a direction.
The computed posterior and the updated support are jointly denoted as $\hat{p}_{(\Theta_{N_{\text{LFI}}})}$. For the MoG posterior, it is implied that the MoG has been parameterised in conjunction with an underlying support. Thus, technically, when sampling from the posterior, we sample within the corresponding support. This makes our algorithmic variants integrable to any framework compatible with the standard BayesSim.
IV-B BayesSim-MODE: mode-shift-based expansion
Algorithm 3 BayesSim-MODE
1:function BayesSim_MODE($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r},\Theta_{0},$...)
2:  Args: (for $N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r}$ see alg. 1);
3:     $\Theta_{0}=[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$: initial support bounds;
4:     $\nu_{\text{TH}}$: mode shift thresh; $\rho$: bounds prox. thresh;
5:     $\tau$: MoG weight sum thresh; $\eta$: expansion rate;
6:     $\Phi$: phys. feasible domain limits
7:  $D\leftarrow\lvert\boldsymbol{\theta}_{\min}\rvert;\;p_{0}\leftarrow\tilde{p};\;i\leftarrow 0;\;\mathcal{D}\leftarrow\{\}$
8:  while $i<N_{\text{LFI}}$ do
9:   $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}\leftarrow$ Simulate NN $\pi_{\boldsymbol{\beta}_{0}}$ rollouts, $\boldsymbol{\theta}\sim p_{i}$
10:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}$ $\triangleright$ Dataset augmentation
11:   Train $q_{\phi}$ over $\mathcal{D}$
12:   $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto(p_{i}(\boldsymbol{\theta})\mathbin{/}\tilde{p}(\boldsymbol{\theta}))q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$
13:   // MODE: Mode-based bounds expansion
14:   // Skip expansion in iteration 0
15:   while $i>0\And d<D$ do $\triangleright$ Iterate param. dims
16:     $r^{(d)}=\lvert\theta_{\max}^{(d)}-\theta_{\min}^{(d)}\rvert \triangleright$ Range size
17:     $W_{L}^{(d)}\leftarrow 0$; $W_{R}^{(d)}\leftarrow 0$
18:     for each Gaussian $k\in$ MoG do
19:      $\nu_{i,k}^{(d)}\leftarrow\mu_{i,k}^{(d)}-\mu_{i-1,k}^{(d)}$ $\triangleright$ Mode shift
20:      // Normalised proximity to bounds
21:      $z_{k}^{(d)}\leftarrow\frac{(\mu_{i,k}^{(d)}-\theta_{\min}^{(d)})}{r^{(d)}}$
22:      // Check mode shift & bound proximity
[…]
24:        $W_{L}^{(d)}\leftarrow W_{L}^{(d)}+w_{i,k}$
[…]
26:        $W_{R}^{(d)}\leftarrow W_{R}^{(d)}+w_{i,k}$
27:      end if
28:     end for
29:     if $W_{L}^{(d)}>\tau\And(\theta_{\min}^{(d)}-\eta r^{(d)})\in\Phi$ then
30:      $\theta_{\min}^{(d)}\leftarrow\theta_{\min}^{(d)}-\eta r^{(d)}$ $\triangleright$ Expand left
31:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\min}^{(d)}\leftarrow\Phi_{\min}^{(d)}$
32:     if $W_{R}^{(d)}>\tau\And(\theta_{\max}^{(d)}+\eta r^{(d)})\in\Phi$ then
33:      $\theta_{\max}^{(d)}\leftarrow\theta_{\max}^{(d)}+\eta r^{(d)}$ $\triangleright$ Expand right
34:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\max}^{(d)}\leftarrow\Phi_{\max}^{(d)}$
35:     $d\leftarrow d+1$
36:   end while
37:   $i\leftarrow i+1$
38:   $\Theta_{i}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ $\triangleright$ Update bounds
39:   $p_{i}\leftarrow\hat{p}_{(\Theta_{i})}$ $\triangleright$ Update reference prior
40:  end while
41:  return $\hat{p}$
42:end function
BayesSim-MODE (Mode-Oriented Domain Expansion) expands the support boundaries based on the position shift in the posterior-mode estimates between consecutive iterations of inference in conjunction with their proximity to these boundaries (alg. 3). The intuition is that high-confidence mode shift toward the boundaries of the current support may indicate that the true parameters lie beyond them. In contrast to EDGE, which relies on a static edge mass, MODE leverages temporal information.
Similarly to EDGE, the standard BayesSim assumptions apply. We also have MODE’s hyperparameters, which are the mode shift threshold $\nu_{\text{TH}}$, the bounds proximity threshold $\rho$, the MoG weights summary threshold $\tau$, the bounds expansion factor $\eta$, and the physically feasible domain limits $\Phi$. The main difference between MODE and EDGE is that we now calculate the shift of the MoG posterior modes $\boldsymbol{\mu}_{i}$ compared to the previous iteration $\boldsymbol{\mu}_{i-1}$ (line 19) along with their normalised proximity to the bounds (line 21). Thus, we do not check for $\Theta$ adaptation at the end of the first inference iteration. In all subsequent iterations, the MoG details of the previous iteration are accessible through the reference prior.
We accumulate the mixture coefficients (weights) of each mode that satisfies the shift and proximity thresholds for the left or right bound (lines 24, LABEL: and 26) of a parameter $d$. The accumulated weights $W_{L}^{(d)}$, $W_{R}^{(d)}$ are used for the expansion criteria (lines 29, LABEL: and 32). If $\tau$ has been exceeded and if an expansion would be within $\Phi$ (lines 29, LABEL: and 32), we expand it by a factor of $\eta$ (lines 30, LABEL: and 33). If $\tau$ has been exceeded, but the expansion would have been beyond the corresponding $\Phi$ limit, then this becomes the new boundary.
MODE, therefore, trades the EDGE’s requirement of configuring the probability mass thresholds with configuring the mode-shift thresholds. Conceptually, MODE dictates that if there is a shift ($\nu_{\text{TH}}$, $\rho$) of high-certainty MoG modes (mixture weights, $\tau$) towards an edge, then we should extend the support in this direction. This rule treats the posterior mode shift as a temporal sign that there is more value in exploring beyond the corresponding $\Theta$ boundary. However, associating the support expansion with the latest mode shift implicitly links the potential for future improvement with the already demonstrated improvement. This is an important assumption, and it can heavily incentivise either exploration or exploitation depending on the mode-shift momentum.
IV-C BayesSim-CENTRE: centring-based adaptive bounds
Algorithm 4 BayesSim-CENTRE
1:function BayesSim_CENTRE($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r},\Theta_{0},$...)
2:  Args: (for $N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r}$ see alg. 1);
3:     $\Theta_{0}=[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$: initial support bounds;
4:     $\Phi$: phys. feasible domain limits
5:  $D\leftarrow\lvert\boldsymbol{\theta}_{\min}\rvert;\;p_{0}\leftarrow\tilde{p};\;i\leftarrow 0;\;\mathcal{D}\leftarrow\{\}$
6:  while $i<N_{\text{LFI}}$ do
7:   $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}\leftarrow$ Simulate NN $\pi_{\boldsymbol{\beta}_{0}}$ rollouts, $\boldsymbol{\theta}\sim p_{i}$
8:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}_{i}$ $\triangleright$ Dataset augmentation
9:   Train $q_{\phi}$ over $\mathcal{D}$
10:   $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})\propto(p_{i}(\boldsymbol{\theta})\mathbin{/}\tilde{p}(\boldsymbol{\theta}))q_{\phi}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$
11:   // CENTRE: Recentre parameter bounds
12:   for $d=1$ to $D$ do $\triangleright$ Iterate param. dims
13:     $r^{(d)}=\lvert\theta_{\max}^{(d)}-\theta_{\min}^{(d)}\rvert \triangleright$ Range size
14:     // Compute new domain centre
15:     $\mu^{(d)}\leftarrow$ $\hat{p}$ modes weighted mean
16:     ${\theta_{\min}^{(d)}}^{\prime}\leftarrow\mu^{(d)}-\frac{r^{(d)}}{2}$ $\triangleright$ $\min$ candidate
17:     ${\theta_{\max}^{(d)}}^{\prime}\leftarrow\mu^{(d)}+\frac{r^{(d)}}{2}$ $\triangleright$ $\max$ candidate
18:     if ${\theta_{\min}^{(d)}}^{\prime}\notin\Phi$ then $\triangleright$ Check if $\min$ feasible
19:      ${\theta_{\min}^{(d)}}^{\prime}\leftarrow\Phi_{\min}^{(d)}$
20:      ${\theta_{\max}^{(d)}}^{\prime}\leftarrow{\theta_{\min}^{(d)}}^{\prime}+r^{(d)}$
21:     end if
22:     if ${\theta_{\max}^{(d)}}^{\prime}\notin\Phi$ then $\triangleright$ Check if $\max$ feasible
23:      ${\theta_{\max}^{(d)}}^{\prime}\leftarrow\Phi_{\max}^{(d)}$
24:      ${\theta_{\min}^{(d)}}^{\prime}\leftarrow{\theta_{\max}^{(d)}}^{\prime}-r^{(d)}$
25:     end if
26:     $\theta_{\min}^{(d)}\leftarrow{\theta_{\min}^{(d)}}^{\prime};\;\theta_{\max}^{(d)}\leftarrow{\theta_{\max}^{(d)}}^{\prime}$ $\triangleright$ Expand $\Theta$
27:   end for
28:   $i\leftarrow i+1$
29:   $\Theta_{i}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ $\triangleright$ Update bounds
30:   $p_{i}\leftarrow\hat{p}_{(\Theta_{i})}$ $\triangleright$ Update reference prior
31:  end while
32:  return $\hat{p}$
33:end function
BayesSim-CENTRE moves the support boundaries such that in each inference iteration the weighted mean of the posterior modes is in the centre of the adapted support (alg. 4). The intuition is that this zero-assumption approach will allow the inference algorithm to adapt the support to better match the current estimate of $\boldsymbol{\theta}$.
Thus, beyond the standard BayesSim assumptions, we now only have physically feasible domain limits $\Phi$. In each inference iteration, following the posterior $\hat{p}$ update (line 10), we compute the weighed mean $\mu^{(d)}$ of the $\hat{p}$ modes for each dimension $d$ of the support (line 15). We use $\mu^{(d)}$ to compute the candidate bounds of the adapted support $\theta_{\min}^{(d)}$$^{\prime}$, $\theta_{\max}^{(d)}$$^{\prime}$ (lines 16, LABEL: and 17). We then evaluate these candidate bounds for their physical feasibility (lines 18, LABEL: and 22), and adjust them accordingly. The result is the adapted support.
CENTRE, therefore, requires less knowledge of LFI performance in a given task, since it does not have any task-specific hyperparameters, e.g. related to probability mass or mode shift. Also, it does a window sliding support adaptation, compared to EDGE and MODE window stretching.
Conceptually, CENTRE dictates that the true parameters may lie within the area defined by the current MoG posterior peaks. This rule treats the positions and weights of the MoG modes as geometric indications of the spatial range of the optimal search space. This resembles a risk minimisation acquisition that is neither aggressively exploratory nor purely exploitative. It may under-react to small high-density shifts, so its behaviour is conservative and centring-oriented.
V Support adaptation for Lotka-Volterra & M/G/1 queue
We experiment with Lotka-Volterra (section III-A) and M/G/1 (section III-B) to address the following questions:
  1. 1.
    Can BayesSim-EDGE, MODE and CENTRE variants infer a higher fidelity posterior while also adapting the posterior’s support for more realistic simulation?
  2. 2.
    What is the impact of each variant on inference performance, and how practical are they to tune?
V-A Sim2Sim parameter inference setup
For parameter inference, we implement and compare the standard BayesSim algorithm, as originally proposed by [19] (alg. 1), and our heuristic variants (section IV). The hyperparameters of EDGE and MODE are available in App. C. CENTRE, as explained, does not have any hyperparameters.
In all experiments, we use an MDNN for the CDF approximation; its hyperparameters are given in App. B, Table II. We perform $10$ inference iterations for Lotka-Volterra and $15$ for M/G/1 (alg. 1, line 8). Each iteration augments the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled from the latest posterior.
V-B Results and discussion
Figure 7: Lotka-Volterra Sim2Sim inference exponential of negative loss for trajectories (top) and parameters $\boldsymbol{\theta}$ (bottom grid) over $10$ iterations.
Figure 8: M/G/1 Sim2Sim inference exponential of negative loss for trajectories (top) and parameters $\boldsymbol{\theta}$ (bottom grid) over $15$ iterations.
Figures 7, LABEL: and 8 show the Sim2Sim results for Lotka-Volterra and M/G/1 for different configurations of the inference problem. For the Lotka-Volterra support adaptation experiments, to maintain a satisfactory data efficiency rate, we consider the physically feasible domain limits $\Phi=[-6,4]\times 4$, which is the broad support defined in section III-A. Although physical feasibility is not an issue in the M/G/1 experiments, we consider $\Phi:[0,20]\times[0,20]\times[0,0.5]$.
For Lotka-Volterra (fig. 7), we see that with a good support (OK), the inference result consistently approximates the ground-truth $X,Y$ population trajectories. CENTRE and EDGE perform consistently better than using broad support. Although CENTRE is better in terms of trajectory proximity to ground truth, the per-parameter plots show that EDGE is more robust in per-parameter prediction accuracy. MODE underperforms, with its temporal shift heuristic being difficult to tune, especially given the task’s stochastic dynamics. Thus, it is almost as inaccurate as the misspecified support inference. These results also motivate a more thorough exploration of how each physical parameter’s value influences the resulting population trajectories. Qualitatively, our results suggest that $\theta_{2}$ (a predator dies) and $\theta_{4}$ (a prey dies) inference performance is better correlated with ground truth trajectory approximation, followed by $\theta_{3}$ (a prey is born).
For M/G/1 (fig. 8), it is qualitatively evident that its increased stochasticity poses a bigger challenge for LFI. Although MODE performs better overall in per-parameter inference, EDGE is more consistent among iterations, which is evident in the proximity of the calibrated trajectory to ground truth. We also see the adversarial effect of a broad support mainly for $\theta_{1}$ and $\theta_{2}$ inference. $\theta_{3}$ seems to be inferred more robustly overall, possibly because in the chosen domain $\theta_{3}$ perturbations have a smaller impact on system trajectory. We also see that among all the M/G/1 $\boldsymbol{\theta}$ components, the $\theta_{1}$ inference generally has a stronger qualitative correlation with the overall trajectory approximation.
Figures 7, LABEL: and 8 show the exponential of the negative of all resulting loss ($\mathcal{L}$) values ($\exp(-\alpha\mathcal{L})$, with $\alpha=1.0$), compressing higher values and extending lower ones (as $\mathcal{L}\rightarrow 0$). This gives a fine-grained insight into performance differences among our heuristic support adaptation methods. The top subplot of each figure shows the transformed loss of the simulated trajectory to the ground-truth trajectory, as averaged after drawing 2020 $\boldsymbol{\theta}$ samples from each iteration’s posterior and simulating the respective trajectories. Original loss is measured as the Euclidean distance. The respective bottom grids show the transformed losses of the inferred physical parameter values to the ground-truth, as averaged for 2020 $\boldsymbol{\theta}$ samples from each iteration’s posterior. Original loss is measured as the absolute difference.
Overall, EDGE success shows that a probability mass heuristic is a more robust approach to support adaptation in stochastic dynamical systems. However, it needs careful hyperparameter tuning, which can come from prior experience with a task. This prior experience can be collected through manual experimentation and pilot runs. Further delving into this topic [52, 53] is beyond the scope of our work. On the other heuristic variants, CENTRE seems more useful than MODE, which is reasonable considering the intricacy of tuning the mode-shift tracking hyperparameters for a stochastic iterative inference process. However, despite its zero assumptions, CENTRE can also struggle in certain tasks. This can be attributed to the Bayesian LFI’s inherent tendency to place probability mass on the most likely candidates of the given parameter space, which can negatively impact a centring-based sliding window approach.
VI Real2Sim2Real for visuomotor DLO whipping
Figure 9: Overview of the Real2Sim2Real workflow (right) for visuomotor DLO whipping (left). We perform LFI for the posterior distribution $\hat{p}$ over system parameters (Real2Sim). We use $\hat{p}$ to perform domain randomisation while training a PPO agent to perform a DLO whipping task. We deploy and evaluate our sim-trained policy in the real world (Sim2Real).
Following our inference results for Lotka-Volterra and M/G/1, we use BayesSim-EDGE to study a higher-dim dynamical system. This is the prefaced visuomotor DLO whipping task with a sparse outcome-dependent reward that implicitly guides the entire DLO body [54] toward a stack of cubes. We use the integrated Real2Sim2Real framework for visuomotor DLO manipulation presented in [46] (fig. 9).
Algorithm 5 Integrated Real2Sim2Real for DLO whipping
1:Given: $N_{\text{LFI}}$: inference iterations;
2:    $\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}$: initial parameter bounds
3:Assume uniform proposal prior $\tilde{p}(\boldsymbol{\theta})\approx\mathit{U}[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$
4:Assign reference prior $p_{0}\leftarrow\tilde{p}$
5:Train initial policy $\pi_{\boldsymbol{\beta}_{0}}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, $\boldsymbol{\theta}\sim p_{0}$
6:Run 11 $\pi_{\boldsymbol{\beta}_{0}}$ rollout in the real env to collect $\mathbf{x}^{r}$
7:// 1. Real2Sim DLO parameter inference (LFI)
8:if Heuristic Support Adaptation then
9:  $\Theta_{0}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$
10:  BAYESSIM_ADAPT $\leftarrow$ {EDGE, MODE, CENTRE}
11:  // Run alg. 2, or alg. 3, or alg. 4
12:  $\hat{p}\leftarrow$ BAYESSIM_ADAPT($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r},\Theta_{0},$...)
13:else
14:  // Default
15:  $\hat{p}\leftarrow$ BayesSim($N_{\text{LFI}},\tilde{p},\pi_{\boldsymbol{\beta}_{0}},\mathbf{x}^{r}$)
16:end if
17:// 2. Policy training in sim
18:Train policy $\pi_{\boldsymbol{\beta}_{1}}(\mathbf{a}_{t}\mid\mathbf{s}_{t})$, $\boldsymbol{\theta}\sim\hat{p}$
19:// 3. Sim2Real policy deployment
20:Evaluate $\pi_{\boldsymbol{\beta}_{1}}$ in the real env by running 11 $\pi_{\boldsymbol{\beta}_{1}}$ rollout
VI-A Task overview
VI-A1 Initialisation
A robot arm picks up the DLO from a designated position $\mathbf{x}_{0}$ on the table by grasping it near one of its tips and raising it to a designated height $h_{0}$. Both $\mathbf{x}_{0}$ and $h_{0}$, as well as the initial DLO pose, remain fixed for all the simulation and real experiments. The execution of the whipping policy begins once the object has been picked up and raised to $h_{0}$. Our task has dynamic features due to the underactuated manner in which we control the DLO, which dangles from one of its tips, and the momentum forces (drag, inertia) acting on its body once we raise it from the table.
VI-A2 Visuomotor elements
We use two dedicated keypoint clusters to track the DLO and the stack of cubes in the 2D pixel space of an RGB image. In both parameter inference (real2sim) and policy training and deployment (sim2real) experiments, we position-control the robot arm end effector (EEF) by commanding its Cartesian pose. Our simulator uses the IsaacGym attractors implementation, whereas in the real world we use a Cartesian impedance controller [55].
VI-B Simulation setup
We setup our simulation in IsaacGym [56]. The simulated environment contains a Franka Emika Panda 7-DoF robot arm with a parallel gripper on a tabletop. In this workspace, a blue DLO also exists, implemented as a tetrahedral grid, and simulated using the corotational finite element method of the FleX physics engine. The whipping targets are a stack of $6$ distinctly coloured cubes: [[red, green, blue, orange, purple, yellow]], in bottom-up order, the yellow cube at the top being the ideal target for maximum reward.
Consistent with our support adaptation methodology (section IV), in all inference experiments, we consider a support $\Theta_{0}$, from which we initially sample the system parameters, and a broader domain $\Phi$ which is the physical feasibility limit of the support adaptation. Naturally, $\Theta_{0}\subseteq\Phi$, and similarly each inference iteration’s (index $i$) support is $\Theta_{i}\subseteq\Phi$. When training a policy with uniform DR ($\text{PPO-}\mathit{U}$ in section VII-C), we sample over $\Theta_{0}$. When training a policy using an inferred MoG for DR, we sample the support in which we converged by the last Real2Sim inference iteration (alg. 5, line 18).
[…]
VI-C Real-world setup
Our real-world setup closely replicates the simulation.
VI-C1 Camera details
We collect visual observations with a RealSense D435i camera that captures $60$ fps, mounted to view the workspace from the right side. To avoid a more elaborate sim and real camera calibration, which would compromise the portability of our method [57, 58], we position the real camera so that the captured images qualitatively approximate the respective sim images similar to [20, 2].
VI-C2 DLO details
For our experiments, we manufacture $4$ blue DLOs using Shore hardness A-40 (DLO-0 with len. $200\,$\mathrm{mm}$\$), 00-20 (DLO-1 with len. $200\,$\mathrm{mm}$\$ and DLO-3 with len. $290\,$\mathrm{mm}$\$) and 00-50 (DLO-2 with len. $270\,$\mathrm{mm}$\$) silicone polymers [23]. As in simulation, our real DLOs are shaped as grids, with negligible height and width at $15\,$\mathrm{mm}$\$ each. Table I summarises our DLO indexes, parameterisations, relative hardness descriptions, and mass. Items are sorted primarily on increasing length and then on increasing softness (as medium soft $\rightarrow$ soft $\rightarrow$ extra soft). We will be referencing DLOs using their index, or as $\langle\text{Length}\rangle;\langle\text{Shore Hardness}\rangle$”.
VI-C3 Cube details
The real cubes are made of lightweight foam, weigh $1.54\,$\mathrm{g}$\$ and have side length $\approx 25\,$\mathrm{mm}$\$. Their colours and stacking order are the same as in simulation (section VI-B). Since DLO-cube physical interactions are momentary, they have some impact on rollout outcome and can be a reality gap factor, which motivates our choice of uniform DR for cube properties during policy training (section VI-B).
TABLE I: Real DLO indexes and parameterisations
DLO idx. 0 1 2 3
Length ($\mathrm{mm}$) $200$ $200$ $270$ $290$
Shore A-40 00-20 00-50 00-20
Hardness (medium soft) (extra soft) (soft) (extra soft)
Mass ($\mathrm{g}$) $47.47$ $42.94$ $57.8$ $63.94$
VI-D Visual perception setup
Figure 10: Overview of our policy rollout and state perception workflow. In each policy step we infer the DLO and stacked cubes keypoints from segmentation images and use them as state input for the MDNN. Margin colours indicate association between sample images and algorithm parts.
VI-D1 Segmentation images
For efficiency, we focus our observations on the blue DLO and the coloured stacked cubes using segmentation images (fig. 10). For this, we fine-tune the segmentation task version of YOLOv11 [59]. We use a dataset of $311$ manually labelled images featuring real and simulated DLO physical interactions, $128$ of which are dedicated to our DLO whipping task. We apply pre-processing through auto-orientation and resizing to $256\times 256$, and augmentation through random Gaussian blur and salt-and-pepper noise, for a final dataset totalling $745$ images. With this dataset, we train the open source YOLO weights for $64$ more epochs.
VI-D2 Keypoints
We further reduce the dimensionality of our visual observations by using keypoints to track the environment objects of interest, i.e. the controlled DLO and the stacked cubes. For the DLO, we use its detection mask to create a DLO-only segmentation image in which we infer $4$ keypoints in real time. For this, we train a transporter model [4], as implemented by [1], using a dataset of $1500$ random policy rollouts of a simulated visuomotor DLO reaching task [46], while sampling $\boldsymbol{\theta}$ from a uniform prior. In this dataset, we also include a small number of real-world visuomotor DLO reaching policy rollouts, and we train for $50$ epochs. For the $6$ stacked cubes, we treat the centroids of their $6$ individual segmentation masks as their keypoints. These $10$ 2D keypoints are our visual observation vectors.
Tracking a DLO from keypoints that are inferred per timeframe is prone to perception noise [2, 46]. The keypoints may be temporally consistent, but permutation invariance is not guaranteed, and there is pixel position noise. The impact of such stochastic factors can be mitigated with, e.g., kernel mean embeddings [2], but we empirically find that they are not needed in our whipping task, which is corroborated by our Real2Sim inference results (section VII-A). This is due to the lack of extreme deformations because of limited DLO-environment contact. Our task formulation preconditions all trained agents to a baseline behaviour, which limits drag on the table and induces almost momentary contact with the cube stack (section VII-C). Thus, we can use the standard BayesSim implementation with MDNN [19].
VI-E Proprioception & control setup
VI-E1 Observation & action vectors
We control the EEF through Cartesian pose commands (77D vectors). We constrain the EEF motion in 22D, moving only along the $x$ and $z$ axes by controlling the respective deltas. Thus, our policy actions are 22D $\langle dx,dz\rangle$ vectors that we sample in the $0.06,0.06]\,$\mathrm{m}$\$ range to maintain smooth EEF transitions. Our 2828D observations are constructed by concatenating the EEF’s $\langle x,z\rangle$ 22D position (proprioception) and the 2626D visual observation, which is constructed by flattening the $10\times 2$D vector of the $4$ DLO and $6$ cube keypoints, and concatenating it with the 66D vector of flags $\mathbf{k}$ which tracks whether a cube has been knocked off the stack during the ongoing episode. $\mathbf{k}$ follows the default bottom-up order of cubes (section VI-B).
VI-E2 Safety limits
We restrict the EEF motion within the $\langle x,y,z\rangle\in\langle[0.275,0.6],[-0.1,0.1],[0.1,0.5]\rangle\,$\mathrm{m}$\$ world frame coordinates, with the robot arm based on $(0,0,0)$. Whenever the EEF leaves this designated workspace, the episode ends as a failure with a reward of 1-1.
VI-F Real2Sim parameter inference setup
We define a physical parameter vector $\boldsymbol{\theta}=\langle l,E\rangle$, where $l$ is the length of the DLO and $E$ is its Young’s modulus. We want to infer a joint posterior $\hat{p}(\boldsymbol{\theta})$, which contains information on both the DLO’s size and its material properties.
Following alg. 5, we begin by assuming a uniform proposal prior $\tilde{p}(\boldsymbol{\theta})$, which we use to initialise the reference prior $p_{0}=\tilde{p}$. We then perform domain randomisation with $\boldsymbol{\theta}\sim p_{0}$, while training in simulation an initial policy $\pi_{\boldsymbol{\beta}_{0}}$ for our task. We perform a rollout of $\pi_{\boldsymbol{\beta}_{0}}$ in the real environment to collect a reference trajectory $\mathbf{x}^{r}$ with a specific DLO. We then perform iterative LFI using BayesSim. In each inference iteration $i$, we use $\pi_{\boldsymbol{\beta}_{0}}$ as a data collection policy, running $N$ rollouts in simulation to collect a dataset $\{(\boldsymbol{\theta},\mathbf{x}^{s})\}^{N}$, with $\boldsymbol{\theta}\sim p_{i}$, on which we train the CDF approximation $q_{\phi}$. We use $q_{\phi}$ and $\mathbf{x}^{r}$ to compute the posterior $\hat{p}(\boldsymbol{\theta}\mid\mathbf{x}=\mathbf{x}^{r})$. We then update the reference prior $p_{i}=\hat{p}$ and loop again.
We use the standard BayesSim algorithm [19] (alg. 1), and our EDGE variant (alg. 2), whose hyperparameters are given in App. D, Table V. In both cases, we use an MDNN for the CDF approximation. Our MDNN hyperparameters are given in App. B, Table II. As in alg. 1, line 8, we approximate the posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each augmenting the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
For all simulated and real trajectories $\mathbf{x}=\langle\mathbf{S},\mathbf{A}\rangle$, with visual and proprioceptive states $\mathbf{S}=\{\mathbf{s}^{t}\}_{t=1}^{T}$ and applied actions $\mathbf{A}=\{\mathbf{a}^{t}\}_{t=1}^{T}$, we use cross-correlation summary statistics to compute the MDNN input, as:
$\psi(\mathbf{S},\mathbf{A})=(\{\langle\mathbf{S}_{i},\mathbf{A}_{j}\rangle\}_{i=1,j=1}^{D_{s},D_{a}},\mathrm{E}[\mathbf{S}],\mathrm{Var}[\mathbf{S}]),$ (4)
where $D_{s}$ is the state space dim., $D_{a}$ is the action space dim., $\langle\cdot,\cdot\rangle$ denotes the dot product, $\mathrm{E}[\cdot]$ is the expectation (mean) and $\mathrm{Var}[\cdot]$ the variance (std dev) [19]. For readability throughout the text, we interchangeably denote the raw $\mathbf{x}$ and summarised $\psi(\mathbf{S},\mathbf{A})$ trajectories, unless otherwise required.
VI-G Reward design
The lack of DLO keypoint permutation invariance hinders the creation of reward functions specific to DLO parts. Not having a specific tip keypoint [45, 7] to guide toward the cubes makes it challenging to learn a generalist policy, however, it increases the practicality of learning specialist policies for specific distributions of DLO parameterisations through an integrated Real2Sim2Real treatment. We implicitly condition our policies for whole-body guidance by designing a sparse reward function that evaluates only the rollout outcome, i.e., which cubes were knocked down. Thus, instead of tracking a specific tip keypoint, we use the inferred posteriors to implicitly learn successful behaviours using less elaborate observation and reward functions.
VI-G1 Reward function
In each state $s_{t}$, we assume a base reward $r_{t}=0.0$ and check whether the top (yellow) cube has been knocked off, in which case we assign an initial reward $r_{t}=2.0$. We then inspect the stack in reverse order (top to bottom, excluding the top for which we have just checked), and for every knocked cube with index $i$ we apply an exponential reward decay, as $r_{t}\mathrel{-}=0.05\times 2^{4-i}$. Thus, for [[red, green, blue, orange, purple]] (excl. top yellow) cube indexes iterated in reverse as [[4, 3, 2, 1, 0]] we can have the respective penalties [[$0.05$, $0.1$, $0.2$, $0.4$, $0.8$]] applied in $r_{t}$.
This sparse outcome-dependent reward means that the greater the impact on the stack beyond the yellow cube, the greater the penalty to the maximum achievable reward for the episode. The episode ends if the top cube has been hit. A cube is hit/whipped/knocked off if its keypoint’s pixel-space displacement is >0.06>0.06. We compute a keypoint’s displacement as the Euclidean distance (L2 norm) between its current and initial positions.
VI-H Policy learning & Sim2Real deployment setup
For policy learning, we use the Stable Baselines3 [60] PPO implementation, similarly to [40, 43, 42]. We keep the default implementation hyperparameters, which reflect the hyperparameters originally proposed in the seminal work [61]. We train for $60,000$ total steps, with a maximum duration of episodes of $16$ steps, and a batch size of also $16$.
To deploy our PPO policies in the real world, we extend an open-source sample-efficient robotic RL framework [55]. We empirically tune the damping and stiffness of the real impedance controller so that $16$ real-world steps approximate the respective $16$ simulation steps.
VI-I Domain randomisation setup
In popular robotics simulators, such as IsaacGym, re-parameterising an existing deformable object simulation to change physical properties such as stiffness practically requires re-initialising the whole simulation [46]. This makes it difficult to integrate such simulators in an RL environment that follows the well-established gym style [62].
We implement DR by training our RL policies in vectorised environments [60] and launching parallel instances of the simulation, each with a different $\boldsymbol{\theta}$, sampled by the current domain distribution $p(\boldsymbol{\theta})$ [46]. $p$ is either the default uniform distribution $\mathit{U}$, or an inferred MoG posterior (alg. 5, line 18). For more Sim2Real robustness and to minimise the need for camera calibration, we introduce a small randomness in the simulated camera and stack placement position. For the sim camera, we uniformly sample $(x,y,z)$ offsets in $(\pm 0.025,\,\pm 0.025,\,\pm 0.025)\,$\mathrm{m}$\$, and for the stack position, we uniformly sample $\pm 0.02\,$\mathrm{m}$\$, $\pm 5\,$\mathrm{mm}$\$ $x$ and $y$ axis offsets.
We launch $12$ concurrent environments to manage the computational demands of deformable object simulations. This raises the importance of the domain distributions’ descriptiveness, since MoG inaccuracy and imprecision can have worse consequences in small-data experiments. The MoGs are sampled in a low-variance method, thus each component is likely to contribute to the set of domain samples. For all nonadapted support experiments, we sample assuming the default support $\Theta_{0}$, whereas for all adapted support experiments, we sample assuming the broader physically feasible $\Phi$ (section VI-B). To ensure that the samples are within the effective support, we clip outlier values. This adheres to the $\Phi$ treatment during iterative inference as formulated in section IV.
VII Real2Sim2Real DLO whipping performance
We address the following questions:
  1. 1.
    Can BayesSim-EDGE infer a higher fidelity posterior and adapt the posterior’s support for more realistic simulation in a visuomotor DLO whipping task?
  2. 2.
    What is the overall impact of the resulting posteriors in Real2Sim2Real for this dynamic DLO manipulation task, when cross-evaluating over a parametric set of (real) DLOs? Specifically, we explore the impact on:
    1. (a)
      Classification granularity of different DLO $\boldsymbol{\theta}$.
    2. (b)
      Object-centric agent adaptation.
    3. (c)
      Perceived challenge of each DLO to our agents.
Figure 11: Progression of prior samples (left) and their respective MoG posteriors (right, see alg. 1, line 10 and alg. 2, line 10) along $15$ inference iterations for DLO-0, DLO-1, DLO-2 and DLO-3. On prior samples scatterplots, we see the progression of each inference iteration’s samples. Colourmapping indicates the iteration each sample was drawn, the lighter the colour, the later the iteration. Each iteration samples’ bounding box is plotted in dashed lines of the same colour. The bigger circular points indicate the accumulated dataset’s mean per inference iteration. A black dashed line marks its trajectory. We plot and annotate every second accumulated mean. On MoG heatmaps, component means are displayed in blue crosses and colourbars quantify likelihood. MoG progression is annotated with arrows pointing to the position of each iteration’s weighted component mean. We annotate every second weighted component mean. For coherence the heatmaps show only the result of the last iterations (also in fig. 12).
Figure 12: Inferred MoG posterior heatmaps and the domain samples drawn when each MoG is used for DR. MoG component means are shown in blue crosses and colourbars quantify likelihood. On the top, we see the results after $15$ BayesSim iterations, where we sample from a prior with a constant, potentially misspecified, support. Orange arrows indicate the need for support adaptation (DLO-0 and DLO-3). On the bottom, we see the results after $15$ BayesSim-EDGE iterations, where we heuristically adapt the support over inference iterations. Blue arrows annotate these adaptations.
VII-A Support adaptation impact on LFI
Figure 11 shows our Real2Sim results using 2D MoG posterior heatmaps, together with the respective scatter plots of the domain samples used in each inference iteration. We see how the relative increase in samples granularity follows the location of the current belief on the system parameterisation. Each MoG has $4$ components, whose mean, variance, and mixture coefficient are parameterised during inference. The tightness and spread of the posteriors is a qualitative indication of the precision of the inference. The means of different Gaussians capture alternative hypotheses of the reference DLO’s parameterisation. Figure 12 shows only the final MoGs, annotating the need to adapt certain domain boundaries and the respective adaptations.
The MoG variance along a parameter’s dimension and the respective spread of the component means (fig. 12) visually indicate any uncertainty in parameter estimation. In general, we have a more precise inference of softness (Young’s modulus), with length being a more common source of uncertainty, which is reasonable, since we always use $4$ keypoints to track the DLO.
When it comes to support adaptation, we see that for the shorter and stiffer DLO-0 the standard BayesSim inference result strongly indicates that its length and softness values are potentially beyond the lower and upper bounds of their respective ranges. We see how the EDGE probability mass accumulation heuristic leads to adaptation of the respective domain ranges toward more likely parameterisations. Similarly, we see that for the longer and softer DLO-3, although standard BayesSim confidently places its softness within the default domain, there are indications that the length may be beyond the default range limits. Again, we see how EDGE expands the upper bound of the length support just enough to more confidently infer its value.
For the longer DLOs-$\{2,3\}$, we see alternative hypotheses (component means) for $\boldsymbol{\theta}$ being spread mainly along the length axis. For DLO-3 this spread is consistent for both standard BayesSim and EDGE, while for DLO-2 it emerges only for EDGE. Although the discrepancy is relatively small, it is a reminder of the probabilistic nature of our methods in combination with the inherent technical challenge of performing Real2Sim inference in a dynamic DLO manipulation task using only vision and proprioception.
Figure 13: Pixel-space (faded lines) and statistical (bolder lines) reality gap along the $15$ inference iterations using default and adapted support of fig. 11. We average over the timesteps of the simulated and real-world keypoint trajectories and compute the bidirectional Chamfer distances (pixel-space) and cross-correlation summary statistics Euclidean distances. Errorbars show the standard deviation for each iteration. For clarity, we average and report the distances of each iteration’s $25\%$ best-scoring samples.
Figure 13 shows the reality gap of the domain samples collected during inference iterations (fig. 11). We measure this gap for the simulated and real DLO keypoint observations both in the pixel space and in the summary statistics, using the keypoint pixels’ bidirectional Chamfer distances and the cross-correlation (section VI-F) Euclidean distances, respectively. Among both default and adapted support experiments, we observe the greatest convergence for the stiffer DLO-0 and the longer and softer DLO-3. This shows that in our parametric set of DLOs, DLOs-$\{0,3\}$ have the highest deviation from the median $\mu$ of the parameterisation domain. In contrast, there is little extrinsically measurable difference between the standard BayesSim and EDGE’s inference iterations.
These results highlight the need of an end-to-end treatment of the LFI evaluation in dynamic DLO manipulation, in order to show the actual impact of each object-centric inferred posterior. Furthermore, the difference in the informativeness of the statistical gap and the pixel-space gap curves shows the importance of evaluating inference performance on the MDNN’s input space and not just defaulting to the feature space of the high-level task formulation.
VII-B Uncertainty over parameter estimation is reflected in DR
Figure 12 shows the scatter plots of the $12$ domain samples drawn from each MoG, which are used for object-centric policy training. We see how the certainty (sharpness) of a posterior along a dimension results in proportionately tightly clustered domain samples. In addition, the spread of the domain samples shows that some of the alternative hypotheses for $\boldsymbol{\theta}$ contribute to the low-variance sampling, due to the relative magnitude of their mixture coefficients. This is most evident in the softer DLOs-$\{1,3\}$ results for both standard BayesSim and EDGE, and in the expanded DLO-0 support of EDGE. For the expanded DLO-0 support, in particular, the spread can be an indication that we could have iterated further. For EDGE, we also observe that the alternative hypotheses for DLOs-$\{2,3\}$ compromise for in-between clusters of samples, again due to the low-variance sampling and the respective mixture coefficients.
In general, adapted support experiments result in cleaner clusters of domain samples. Although we cannot yet tell how useful these new clusters will be in policy training, this shows that LFI benefits from a reasonably broader support.
VII-C Real2Sim2Real for object-centric agent adaptation
Figure 14: Learning curves of PPO agent training when performing DR using different domain distributions, producing the fig. 12 domain samples. On the top, we have the agents using the default, potentially misspecified, support. On the bottom, we have the agents sampling from domain distributions with a support that has been heuristically adapted using BayesSim-EDGE.
We train (fig. 14) $4$ PPO policies by domain randomising using the $4$ inferred MoG posteriors with the default support $\Theta_{0}$ and $4$ PPO policies by domain randomising using the $4$ inferred MoG posteriors with a potentially adapted (BayesSim-EDGE) support (fig. 12). We also train a policy by performing DR over a uniform distribution in $\Theta_{0}$ ($\text{PPO-}\mathit{U}$) and a policy which assumes that the simulated DLO is parameterised according to the median $\mu$ of $\Theta_{0}$ ($\text{PPO-}\mu$). We evaluate these $10$ policies in each of our $4$ real-world DLOs and a simulated median DLO, parameterised by the $\Theta_{0}$ median $\mu$. We repeat each evaluation $4$ times for a total of $160$ sim2real and $40$ sim2sim policy deployments.
Figure 15: EEF trajectories from default (top) and adapted (bottom) support policies deployment (col. 1-6) in real world (rows 1-4 & 6-9; 44 real DLOs) and in simulation (rows 5 & 10; median DLO). We repeat each deployment 44 times and plot each episode’s measured accumulation of commanded EEF translations along the $x$ and $z$ axes along with the respective soft-DTW barycentre (bolder lines). To indicate trajectories’ duration, we annotate every three timesteps from their beginning. Although $\text{PPO-}\mathit{U}$ and $\text{PPO-}\mu$ are default support policies, we plot them on both top and bottom grids for visual consistency. The histograms (col. 7) show all collected episode rewards on each DLO (row), with the last column (col. 8) giving a performance overview of each grid’s agents to each row’s DLO.
For each experiment, fig. 15 shows in faded lines the trajectories resulting from the accumulation of commanded EEF position deltas ($\langle dx,dz\rangle$ actions) over $4$ repetitions. We plot in bold the soft dynamic time warping (soft-DTW) barycentre of the trajectories, which enables averaging and clustering inhomogeneous time series under the DTW geometry [63]. We report reward histograms (col. 7), with bin colours matching trajectory line colours, and the overall performance for each set of agents for a given DLO (col. 8). For each DLO (row), col. 8 lists: (i) the average reward achieved by each policy; (ii) the average reward across all policies (Avg set Perf.), which is the average performance of all our agents on that DLO; (iii) the $50$th percentile of all rewards across agents for the given DLO, which emphasises overall consistency (Cons.); (iv) the DLO difficulty, computed as $1-(\text{Cons.}-r_{\min})/(r_{\max}-r_{\min})$, where $r_{\min}$ and $r_{\max}$ are the global $\min$ and $\max$ rewards across all sim and real evaluations. In this scaling to $[0,1]$, higher values indicate lower robustness relative to the global performance range, and thus a DLO that is more challenging to use. This interpretation of our results views the DLO as the tool used by an agent to physically interact with the environment.
Figure 16: Extrinsic evaluation timelapses of real-world PPO-0 and $0^{\prime}$ policy evaluations using DLO-0, in conjunction with fig. 15. A best-case outcome can result in almost identical trajectories, despite PPO-0 being generally less stable. Dotted white and red lines mark the initial EEF $z$ and table surface respectively. A purple indicator on top left of each image marks the initial EEF position, with dotted purple lines marking its visual displacement.
Figure 17: Extrinsic evaluation timelapses of real-world PPO-$1^{\prime}$ and $3^{\prime}$ policy evaluations using DLO-3, in conjunction with fig. 15. Annotations follow fig. 16. We observe the greater DLO distance from the table on $t\in[3,10]$, annotated with light blue arrows with black border along the PPO-$3^{\prime}$ trajectory.
Consistent with recent work on Real2Sim2Real integration [46], we observe EEF motion patterns that adhere to the properties of the domain distributions used for each agent’s training. These patterns show adaptation of agent performance to the inferred physical parameterisation of the corresponding DLO. An indicative subset of these observations is extrinsically corroborated by figs. 16, LABEL: and 17. All performance observations are referenced in olive annotations. Policy underlines match plot line colours with reference to the DLO used for the domain distribution inference (fig. 14).
From the reward histograms, we see that policies trained using an adapted support generally result in higher rewards. This is most evident for the $\text{EDGE-MoG}_{\text{DLO-0}}$ policy ( $\text{PPO-}0^{\prime}$), compared to the $\text{MoG}_{\text{DLO-0}}$ policy ($\text{PPO-}0$) (1) and also for $\text{EDGE-MoG}_{\text{DLO-2}}$ policy ( $\text{PPO-}2^{\prime}$) compared to $\text{MoG}_{\text{DLO-2}}$ policy ($\text{PPO-}2$) (2). We also see that support adaptation has no impact for DLO-1, for which we have dedicated policies $\text{MoG}_{\text{DLO-1}}$ ($\text{PPO-}1$) and $\text{EDGE-MoG}_{\text{DLO-1}}$ ( $\text{PPO-}1^{\prime}$). This is reasonable, as the very short and soft DLO-1 aligns well with the default support $\Theta_{0}$, and thus should need the least support adaptation of all our DLOs.
For the longer and softer DLO-3, with its dedicated policies $\text{MoG}_{\text{DLO-3}}$ ($\text{PPO-}3$) and $\text{EDGE-MoG}_{\text{DLO-3}}$ ( $\text{PPO-}3^{\prime}$), we see that support adaptation leads to a significant deployment performance gain when $\text{PPO-}3^{\prime}$ controls DLO-2. For PPO-$3^{\prime}$ controlling DLO-$3$, although the motion is useful, it is not fast enough for the impact on the top cubes to be properly recorded in $16$ timesteps (fig. 17, bottom). We see that the $\text{PPO-}1^{\prime}$ policy, which is preconditioned for the equally soft but shorter DLO-1, appears to perform better for DLO-3 by achieving the optimal reward of $2.0$ in $2/4$ rollouts (3). However, extrinsic evaluation shows that the behaviour is useful in just $1/4$ rollouts. On the other hand, while $\text{PPO-}3$ had some utility for DLO-1, the over-adapted $\text{PPO-}3^{\prime}$ does not, having specialised in longer bodies.
This divergence of expected and actual $\text{PPO-}3^{\prime}$ performance shows that, despite the aforementioned benefits, combining support adaptation and zero-shot deployment for dexterous visuomotor control tasks, such as DLO whipping, can still leave us exposed to an increased reality gap. However, it also shows how the perceived parameterisation of a DLO, as reflected in a domain distribution, can influence the skill exhibited by the corresponding agent in using various (real) DLOs as tools for physical interactions (cube whipping).
The overall scores give further insight into the impact of support adaptation on object-centric agent performance. In the default support MoG posteriors of fig. 12, we see that DLO-0 needs an adaptation of at least the stiffness subdomain. BayesSim-EDGE achieves this, along with some proportionately smaller length subdomain adaptation, leading to increased performance for the corresponding PPO-$0^{\prime}$ policy. In contrast, DLO-$1$ does not need any support adaptation, but BayesSim-EDGE results in a sharper posterior, leading to an approximately median performance for PPO-$1^{\prime}$. The MoG inferred by BayesSim-EDGE for DLO-$2$ leads to PPO-$2^{\prime}$, which overall is more useful than PPO-$2$. In the DLO-$3$ EDGE posterior, we see a length adaptation. However, the latent interaction of the length subdomain adaptation with the Young’s and density subdomains results in PPO-$3^{\prime}$ being even more useful for DLO-$2$ than DLO-$3$.
Our agents sim2sim performance on DLO-$\mu$ corroborates our previous median performance observation, as we see the policies that are preconditioned through DR on DLOs-$\{1,2\}$ being the most robust, along with $\text{PPO-}\mathit{U}$. This is closely related to our previous discussion (section VII-A) on the relative proximity of DLOs-$\{1,2\}$ to the parameter space median $\mu$.
Extrinsically, we observe that the policies of the adapted support DLO-$\{0,2,3\}$ posteriors, i.e. PPO-$\{0^{\prime},2^{\prime},3^{\prime}\}$, show a less varied performance among rollouts for their respective DLOs (4-6). For DLO-0, this is evident in the shorter soft-DTW trajectory. In contrast, we see that for DLO-1 the rollout variance of $\text{PPO-}1^{\prime}$ increases and its DLO-1 performance decreases compared to $\text{PPO-}1$. However, we also see $\text{PPO-}1^{\prime}$ being useful for longer DLOs-$\{2,3\}$, overall being competitive to $\text{PPO-}2^{\prime}$ performance. Another, more subtle, pattern is that PPO-$\{0^{\prime},1^{\prime},2^{\prime},3^{\prime}\}$ generally adapts better than PPO-$\{0,1,2,3\}$ to the manipulated DLO length, as indicated by the EEF distance from the table surface during movement (7). Moving closer to the table risks dragging the DLO on its surface. This physical interaction is hard to formulate in RL, but can nevertheless have adversarial consequences. We mitigate them by using an appropriate distribution and support for DR, which helps us avoid a more complex problem formulation.
In general, we see that support adaptation is a useful mechanism that can increase object-centric agent performance. Still, it should be used carefully in cases where it may be redundant. This is important in iterative LFI, since the Bayesian tendency to attribute the inference target to the closest samples drawn within an effective support (section II-D) can increase the impact of false positives, from more uncertain earlier inference iterations, on support adaptation. This can be mitigated with careful hyperparameter tuning.
VIII Conclusion
In LFI, we usually assume an arbitrary support for the prior. However, this exposes us to a potential misspecification issue. In an integrated Real2Sim2Real framework for robotics, this issue can be propagated to policy learning and deployment. We treat the misspecified support issue as an information acquisition problem and explore how heuristics can guide support adaptation in iterative LFI, in order to converge to a more useful support in conjunction with inferring the posterior $\hat{p}(\boldsymbol{\theta})$.
For our first two contributions, we empirically expose the above issue and propose three heuristic variants of the BayesSim LFI method [19] (EDGE, MODE, and CENTRE). Each variant heuristically monitors posterior features, such as mass accumulation and mode shift among inference iterations, and adapts the support accordingly. We evaluate our variants using two stochastic dynamical benchmarks, the Lotka-Volterra and M/G/1 queue, for Sim2Sim experiments in which we need to recover from a misspecified support. The EDGE probability mass heuristic performs most robustly, with the caveat of requiring careful hyperparameter tuning.
For our third and fourth contribution, we design a visuomotor DLO whipping task, a higher-dim stochastic dynamical system, to study the impact of support adaptation within an integrated Real2Sim2Real framework. We use our EDGE heuristic to adapt the support among Real2Sim inference iterations, successfully augmenting BayesSim’s capacity to infer fine physical parameterisation differences, such as different length and softness amongst visually similar real DLOs.
We then explore the Sim2Real impact of different DLO posteriors, with and without support adaptation. Our results show that support adaptation leads to stronger object-centric agent adaptation and improved overall task performance for $3/4$ real DLOs ($\{0,2,3\}$). However, we also see how the adapted support posteriors can lead to a divergence of expected and actual agent specialisation, e.g. $\text{PPO-}1^{\prime}$ is more useful when deployed on DLOs-$\{2,3\}$ than DLO-1.
Overall, we take a first step towards relaxing LFI’s dependence on a priori assumptions. The limitations of our work motivate the next steps. Although we show how an initial support can be adapted during inference, this requires defining a physically feasible domain $\Phi$ to limit the search. However, defining $\Phi$ requires experience with the given task, which is laborious to acquire manually and difficult to automate. Instead, we could condition support adaptation on simulation quality standards being met. This would interpret increased keypoint noise, or generally observation deviation from the mean, as perception error [25, 26], which can be due to simulation failure, since our perception models are trained on successful trajectories. We could also use more but shorter inference steps, sampling fewer trajectories for MDN training, and swap explicit heuristics with supervision of how the posterior responds to random support expansions.
Appendix A The misspecified support issue for M/G/1
Figure 18: The misspecified support issue for M/G/1. We plot the progression of prior samples (top) and their respective MoG posteriors (bottom) along $10$ inference iterations. Layout, colour coding, and annotations follow Figure 5.
Figure 18 shows how a broader support (3rd col.) hinders the inference of a precise posterior for the M/G/1 task. As discussed in section III-C, M/G/1 is less helpful in exposing the misspecified support issue, as its formulation is inherently forgiving of outlier samples. The misspecification and excessive broadening implications here are mostly evident in the overall posterior confidence (see heatmap values on the colourbars) and consequently in posteriors’ sharpness.
Appendix B Mixture density networks hyperparameters
For all our BayesSim and variants (EDGE, MODE, CENTRE) inference experiments, we approximate the CDF with an MDNN [34], parameterised and trained per Table II.
TABLE II: MDNN hyperparameters
Hyperparam. value
mixture components 44
layers $3\times[1024]$, fully connected
optimiser Adam
learning rate 1e51e-5
batch size $100$ (Lotka-Volterra, M/G/1), $10$ (DLO whip)
Appendix C Sim2Sim experiments hyperparameters
Tables III, LABEL: and IV feature the hyperparameters of BayesSim-EDGE and MODE for the Lotka-Volterra and M/G/1 support adaptation experiments (section V).
TABLE III: BayesSim-EDGE Sim2Sim hyperparameters
Hyperparam. Symbol Lotka-Volterra M/G/1
edge zone fraction $\delta$ $0.1$ $0.2$
prob. mass thresh $\tau$ $0.005$ $0.001$
edge expansion factor $\eta$ $0.2$ $0.2$
TABLE IV: BayesSim-MODE Sim2Sim hyperparameters
Hyperparam. Symbol Lotka-Volterra M/G/1
mode shift thresh $\nu_{\text{TH}}$ $0.01$ $0.01$
bounds prox. thresh $\rho$ $0.4$ $0.4$
MoG weight sum thresh $\tau$ $0.05$ $0.05$
expansion rate $\eta$ $0.2$ $0.2$
TABLE V: BayesSim-EDGE Real2Sim hyperparameters
Hyperparam. Symbol DLO whip
edge zone fraction $\delta$ $0.1$
prob. mass thresh. $\boldsymbol{\tau}$ $\langle 0.003,0.005\rangle$ (DLO-$\{0,1,3\}$),
or $0.005$ (DLO-22)
no significant difference, see section VII-A
edge expansion factor $\eta$ $0.2$
Appendix D Real2Sim experiments hyperparameters
Given $\boldsymbol{\theta}=\langle l,E\rangle$ (section VI-F), Table V presents the hyperparameters of BayesSim-EDGE for the DLO whipping task Real2Sim support adaptation experiments (section VII).